4. Each member of a group of n players rolls a die.
(a) For any pair of players who throw the same number, the group scores 1 point.
Find the mean and variance of the total score of the group.
(b) Find the mean and variance of the total score if any pair of players who throw
the same number scores that number.
5. Every package of some intrinsically dull commodity includes a small and exciting plastic
object. There are k different types of object, and each package is equally likely to
contain any given type. You buy one package each day.
(a) Find the mean number of days which elapse between the acquisitions of the n
th
new type of object and the (n + 1)th new type.
(b) Find the mean number of days which elapse before you have a full set of k objects.
1
Expert's answer
2020-11-17T06:30:09-0500

4)a) Let Si denote the score obtained by the players who throw i,i=1,2,…,6 , and let Xi be the number of people who throw i, and 1ij  be the indicator function which is 1 only if the jth person throws i. So, we have Xi=∑nj=11ij.  E[Si|Xi]=Xi(Xi−1)/2 and so E[Si]=(n2−n)/72 and therefore E[S]=(n2−n)/12, where S=∑6i=1Si is the total score.

b) The indicator variables Yij that are 1 if i and j throw the same number are pairwise independent and can thus be used to find both the expectation and the variance. The expectation of Yij is 1/6 and the variance is "\frac16\cdot\left(1-\frac16\right)=\frac5{36}" ,so the expectation of the score is "\binom n2\cdot\frac16=\frac{n(n-1)}{12}" and the variance of the score is "\binom n2\cdot\frac5{36}=\frac{5n(n-1)}{72}"

5)suppose there are 6 types of objects

the probabilities of getting a new type go on decreasing from 1 to 5/6, to 4/6....as we get more & more types, and using E[x] = 1/p, the expected # of days to get the next type goes on increasing

E[days to get 1st type] = 6/6 ...... 6/(7-1)

E[more days to get 2nd type] = 6/5 ..... 6/(7-2)

E[more days to get 3rd type] = 6/4 ..... 6/(7-3)

a. E[more days to get (j+1)th new object] = c/(c-j)

b. E[days to get full set] = Σ(c-j+1)